I need help solving this problem. Please help.
You need a 21% salt solution to fill your fish tank. How much of the 15% salt solution should be added to 12 gallons of the 25% salt solution to create the desired salt solution?

Answer by ptaylor(2198)

Let x=amount of 15% solution needed
Total solution after the mixture takes place is 12+x gal, right?
Now we know that the amount of pure salt in the 15% solution(0.15x) plus the amount of pure salt in the 12 gal (0.25*12) has to equal the amount of pure salt in the final mixture (0.21*(12+x)). So our equation to solve is:
0.15x+0.25*12=0.21(12+x) get rid of parens and simplify
0.15x+3=2.52+0.21x subtract 0.15x and also 2.52 from both sides
0.15x-0.15x+3-2.52=2.52-2.52+0.21x-0.15x collect like terms
0.48=0.06x divide both sides by 0.06
x=8 gal--------------- amount of 15% solution needed
CK
0.15*8+0.25*12=0.21*20
1.2+3=4.2
4.2=4.2

Hope this helps----ptaylor